OpenStudy (anonymous):

All of the following are equivalent, except _____. (12 + 3)x (12 + 3x)x 15x 12x + 3x

OpenStudy (mathmale):

( ) in this context, preceded or followed by x, indicates required multiplication. Try multiplying out the first two expressions before comparing the results with the second two expressions.

OpenStudy (anonymous):

anything times 1 is itself

OpenStudy (simplymarie_x):

(12 + 3)x = 3x + 12x = 15x 12 + 3x)x = 3x^2 + 12x 15x = 15x 12x + 3x = 15x So out of which of those isn't the same?

OpenStudy (anonymous):

whats the answer to b?

OpenStudy (simplymarie_x):

The answer to b is 3x^2 + 12x. Thats as much as it simplifies to

OpenStudy (anonymous):

oh ok thx